OpenStudy (brittany):

simplify this expression 3 square roots of ninne

OpenStudy (angela210793):

\[3\sqrt{9}\]? or \[\sqrt[3]{9}\]

OpenStudy (anonymous):

\[3\sqrt{9}=3*3=9\]

OpenStudy (anonymous):

do you mean \[\sqrt[3]{9}\] or \[3\sqrt{9}\]

OpenStudy (aravindg):

9

OpenStudy (brittany):

3√9

OpenStudy (anonymous):

or cube root of 9

OpenStudy (aravindg):

3*3=9

OpenStudy (angela210793):

then mathematicalpro is right

OpenStudy (aravindg):

thx